Given an event, A, what equation gives you the probability that event A will NOT occur?
What is ....
P(A') = 1 - P(A)
or
P(A')+P(A)=1
A coin is tossed 3 times. What is the probability that it lands on heads exactly one time?
What is...
0.375
Events that do not impact each other's probability.
What is...
independent
The probability of something always lies between?
What is ...
0 and 1 (inclusive)
When exploring the addition rule and conditional probability we looked to see who could do the ____________ and the Evil Eyebrow.
What is ...
the Taco Tongue?
Given two independent events, A and B, what is the probability that event A and B occurs.
What is...
P(A and B) = P(A) * P(B)
Armando is a high school baseball player. He gets a hit 70% of the time. That means his average is 0.70. What is the probability that he gets his first successful hit on his third swing?
What is ...
0.3*0.3*0.7= 0.063
two events that cannot occur simultaneously
What is....
mutually exclusive or disjoint
Simulated probabilities tend to get closer to the true probability as the number of trials increases.
What is...
the Law of Large Numbers?
The Venn Diagram was often paired with a ___________ table.
What is ...
Two way?
Given two mutually exclusive events, A and B, what is the probability that event A or event B occurs?
What is ...
P(A or B)= P(A) + P(B)
The two equations used to determine independence between 2 events.
What is...
P(A|B)=P(A)
or
P(B|A)=P(B)
two events that cannot occur together, but one must happen
What is...
complementary
A way to model random events, such that the outcomes closely match real world outcomes.
What is ...
a simulation.
Simulating a 50% chance could be used for what real-life situation?
What is...
a True or False test?
Given two events, A and B, what is the probability that event A and B occurs.
What is...
P(A and B) = P(A) * P(B|A)
90% of the people who have a particular disease will have a positive result on a given diagnostic test (90% accuracy). If 5% of a certain population has the disease, what percent of that population would test positive for the disease?
What is ...
14% or 0.14
events that do not impact each other's probability
what is ...
dependent
The probability of a sample space should total up to
What is ...
1.
A way to display or model all the different outcomes of our sample space.
What is....
a Tree diagram?
Given two events, A and B, what is the probability that event A or event B occurs?
What is ...
P(A or B)= P(A) + P(B) - P(A and B)
or
P(A or B)= P(A) + P(B) - P(A)*P(B|A)
Two events are this if they share no outcomes in common
What is ...
disjoint or mutually exclusive
contains all of the possible outcomes
what is...
sample space
Interpret this probability: The probability of flipping a coin and getting heads is 50%.
What is ...
After many, many flips of a coin the percent of times of getting heads will approach 50%.
When outcomes are equally likely we can put the # of outcomes we want over the...
What is ...
the Total # of outcomes in the sample space?